OpenStudy (anonymous):

A container has 50 electronic components, of which 10 are defective. If 6 components are drawn at random from the container, what is the probability that at least 4 are good?

OpenStudy (anonymous):

0.26 0.42 0.75 0.91 1.00

OpenStudy (nurali):

P(defective component) = 1/5.....so, P(good) = 4/5 P(at least 4 good) = P(4 good) + P(5 good) + P(all 6 good) => 6C4(4/5)⁴(1/5)² + 6C5(4/5)⁵(1/5)¹ + 6C6(4/5)⁶(1/5)⁰ => 0.2458 + 0.3932 + 0.2621 i.e. 0.9011